Just Do It Inc. manufactures widgets. The company has the capacity to produce 100,000 widgets per year, but it currently produces and sells 75,000 widgets per year. The following information relates to current production:

Sale price per unit..................................$41
.Variable costs per unit:
..Manufacturing....................................$24
..Marketing and administrative............$5
Total fixed costs:
..Manufacturing.............................$80,000
.. Marketing and administrative...$23,000

If a special sales order is accepted for 8100 widgets at a price of $39 per unit, and fixed costs increase by $12,000, how would operating income be affected? (NOTE: Assume regular sales are not affected by the special order.)

A) Increase by $94,000

B) Decrease by $68,000

C) Increase by $68,000

D) Increase by $108,500

E) Increase by $69,000
Answer #1

Total sales order = 8,100 units

Selling price = $39

Variable cost = 24 + 5 = $29

Contribution margin = 39 - 29 = $10

Total contribution margin = 8,100*10 = $81,000

Fixed cost will increase by $12,000 (Rest all fixed expense will not be included as they are fixed in nature)

Net increase in Operating income

= Contribution margin - Fixed cost

= 81,000 - 12,000

= $69,000

Therefore the correct option is E

If you find the answer helpful please upvote.